Solve by the elimination method
7r-5s=2
5r+ 7s = 86
I really want to see how to solve this one step by step thank you
Laurie

Answer by ankor@dixie-net.com(6693)

solve by the elimination method
:
7r - 5s = 2
5r + 7s = 86
:
What we want to do, is multiply each equation by a value that will make the
coefficients of s the same is each equation:
:
Multiply the 1st equation by 7; and the 2nd equation by 5, results:
49r - 35s = 14
25r + 35s = 430
-----------------Addition eliminates s, so we can solve for r
74r + 0s = 444
74r = 444
r = ; divided both sides by 74
r = 6
:
Find s by substitution of r, in the 1st equation
7(6) - 5s = 2
42 - 5s = 2
-5s = 2 - 42
-5s = -40
5s = 40; multiplied both sides by -1
s =
s = 8
;
;
Check our solutions in the 2nd original equation
5r + 7s = 86
5(6) + 7(8) =
30 + 56 = 86; confirms our answers,
:
:
Did this shed some light on this stuff for you??